# The Hands-Up Problem and How to Deal With It: Secondary School Teachers’ Experiences of Debugging in the Classroom

### v1

@@ 3.3 Analysis
We analysed teachers’ interviews using thematic analysis (TA) (Braun and Clarke, 2006). Braun and Clarke (2021) describe TA as a ‘cluster’ of methods rather than a single analytic procedure. Out of the three ‘schools’ they propose, we describe our analysis as ‘codebook TA’. This school maintains the core values of reflexive TA, in which the subjectivity and experiences of the researcher are key aspects of the analysis, while incorporating some pragmatic approaches that seek reproducible coding (Braun and Clarke, 2021). In other words, we wished to generate themes that identify meaning beyond teachers’ actual responses while ensuring the codebook we developed was consistently applied across the interview data.
In practice, the procedure followed was still guided by Braun and Clark’s original article on TA (Braun and Clarke, 2006). This began with multiple passes of the interview transcripts. The first of these, as previously mentioned, was used both for correcting the automatically generated transcripts and generating initial memos. This was performed by the first author. The second author then read through the corrected transcripts and wrote initial memos.
Teachers’ interview transcripts were then migrated to NVivo 12, which was used for the rest of the analysis. Each transcript was then read through again by the first author to inductively generate open descriptive codes and note further memos. Through further inspection of the interview data, open codes were inductively abstracted into higher-level latent themes and sub-themes that aimed to provide an interpretive, and hence subjective, account of the data. This process of theme generation involved repeatedly restructuring the codebook, redefining the themes and sub-themes, and rewriting the report of the codebook, encompassing steps 4-6 in Braun and Clarke (2006). This process was conducted by the first author with feedback from the second author after some of the iterations. Our approach was hence more iterative than linear, which is considered a normal part of TA (Braun and Clarke, 2006).
@@ 3.4 Reliability and Validity
To strengthen the content validity of the interview structure, the interview questions were iteratively refined by both authors to ensure they appropriately discussed teachers’ experiences, perspectives, and observations specifically related to debugging. This involved a pilot interview with a single teacher to assess the wording of the questions and the timings of the interview. This teacher was personally contacted by the first author and received the same bookshop voucher as the other participating teachers.
The reliability with which the codebook was applied was improved by seeking agreement between the authors without calculating an interrater reliability (IRR) coefficient, which Campbell et al. (2013) call ‘negotiated agreement’. We wanted to improve the consistency with which the first author coded the interview data according to their codebook definitions (also known as intracoder reliability (O’Connor and Joffe, 2020)). Seeking agreement, by identifying points of disagreement, is appropriate for this case (Campbell et al., 2013; McDonald et al., 2019). This was particularly important given the first author was inexperienced at qualitative analysis. Negotiated agreement hence improved the clarity of definitions and the application of the codebook.
Calculating an IRR coefficient was not required for the above process and was considered detrimental for our study. Our interview transcripts were long and conversational, including frequent tangents and interruptions. Naturally, this makes it difficult for coders to agree on units of meaning, which is required for calculating agreement with IRR coefficients. Combined with the complexity of the codebook, any IRR coefficient is likely to be positively skewed and less representative of the reliability of the coding system (Campbell et al., 2013). Furthermore, there is little consensus on acceptable IRR coefficient values for semi-structured interview data, which undoubtedly differs from research using other qualitative data or with different objectives (Krippendorff, 2004; Campbell et al., 2013).
Agreement was hence sought as follows. As described in the previous section, the first author developed the coding frame with regular input from the second author. Two interview transcripts (Gideon and Hannah) were then randomly selected to be coded by the second author, in line with the guidelines suggested by O’Connor and Joffe (2020). They were provided with the codebook, complete with definitions, and an uncoded version of the interview transcripts. Once coded by the second author, disagreements in the coding of sub-themes were identified using the coding stripes feature on NVivo. Disagreements and ambiguous sub-themes were then discussed and where necessary, revisions were made to the structure and definitions of the codebook.
@@ 3.5 Limitations
A limitation of our methodology was that we kept the final themes close to research questions (i.e., closely related to debugging). This was done to avoid generating an overly large codebook, but in the process likely missed some information that contributes to teachers’ experiences of debugging in the classroom. Additionally, our sample was, within certain constraints, self-selecting. Although we would not be seeking a generalisable population for an in-depth qualitative study in the first place, the participating teachers are likely engaged with computing education research to some degree. As a result, they may be more reflective or intentional about their teaching practice than most teachers, potentially affecting the themes we identified.
@@ 4 Results
The complete code definitions are available to view online [^4] https://github.com/LaurieGale10/secondary-teachers-experiences-debugging-study-materials. We now report each of the themes, with sub-themes reported in italics, supported by teachers’ quotes [^5] All teachers have been given pseudonyms and all quotes have had fillers and contractions removed. from the interviews.
@@ 4.1 Barriers to successful debugging
A particularly widespread theme identified in many parts of the interviews was the obstacles that impede students from successfully debugging their programs. These challenges related to schools’ allotted time for teaching programming, students’ attitudes, and the nature of the Python language and environment.
@@
Barriers related to the difficulties of learning to program (programming-related), some of which were specific to Python, also make debugging difficult for students. The difficulty of getting syntax exactly right and conceptual difficulties, sometimes manifesting as logical errors, were both mentioned as struggles for students. Many teachers described the common errors students encounter, mostly related to minor incorrect formatting, illustrating the ‘syntax barrier’ (Denny et al., 2011) present for many beginner programmers. Esther makes the point that such syntax errors may not be due to a lack of correct recall, but a “a lack of the deeper understanding of what the language is actually trying to do”.
@@
Gideon emphasises that students’ negative attitudes span from debugging as a skill to computing as a subject. He also alludes to a lack of success and motivation as drivers for these negative attitudes. The former relates to the ‘performance accomplishment’ determinant of self-efficacy (Bandura, 1978), which is derived from a learner’s initial experiences with a subject. For most students in England, initial experiences of programming come at primary school. However, unhelpful attitudes could also arise from students’ experiences of learning more generally, which Isabel believes to be the case.
@@
The effect of these attitudes often manifests as a lack of motivation to debug their programs. This has consequences for the teacher, who has the challenge of motivating these students.
@@ 4.2 The emotional nature of debugging
As well as attitudes towards debugging, teachers discussed a range of emotional reactions they observe when their students are debugging. These reactions are generally physical or verbal, observed at the point of encountering an error, or after successful or unsuccessful attempts to resolve an error. Visible reactions are not universal, however. Some teachers specifically pointed out the reserved reactions of quieter students.
Most reactions represented negative emotion during debugging struggle, conveying feelings such as frustration and disappointment. An interesting observation mentioned by several teachers is the emotional attachment students develop towards their programs and the distress this can cause when errors occur. Getting an error in students’ own programs can be far more costly than an erroneous program written by the teacher, leading to more dire emotional reactions. Felix describes this as potentially depressing for students.
“When they’re starting to write upwards of ten lines of code, you’re invested in getting it right and then if it fails that can flip and that can be quite depressing.” - Felix
@@
“Once they’ve hit the brick wall, that’s kind of the green light to go, ‘I’m just gonna do anything.’ Sit there in silence, they’re not, they’re not causing any commotion.” - Gideon
Despite the lack of visible frustration, disengagement may indicate more severe emotional distress, as students may not be able to bear any more errors in their programs. Alternatively, students may disengage due to the difficulty of the error. Either way, some students’ motivation to debug is reduced amid repeatedly encountering errors. It can be difficult for busy teachers to notice these students who, as Gideon mentions, are ‘not causing any commotion’.
@@
Felix’s observations allude to the physical enactment of students’ emotions, demonstrating the sheer joy that can be experienced from debugging. However, such positive reactions were rarely mentioned, perhaps due to the perseverance required to experience them.
We also found some teachers alluding to the resilience of motivated students, who do not experience such polarising emotions when debugging. Instead, these students remain composed and try different solutions, not afraid to add further errors to their programs. Hannah explains that most of her students have such resilience.
@@
‘The resilience to bother to debug’, to a large extent, comes from the student. From Hannah’s perspective, this derives from students’ interest in programming. Other factors are undoubtedly at play, such as the debugging-specific teaching strategies teachers use (see Section 4.5), their work in normalising errors in the classroom (see Section 4.6), and their access to computers at home. Hannah teaches in a selective school, where high levels of interest are perhaps not as surprising. Generally, however, negative emotions and disengagement were the most widely discussed ideas we identified in this theme.
@@ 4.3 Reliance on the teacher
A combination of the barriers to successful debugging and its emotional nature contribute to many students relying on the teacher for debugging support. This is typically embodied by students raising their hands when encountering errors. Often, there are enough hands up to make it unrealistic for the teacher to effectively help every student requesting support.
Several teachers, such as Chloe and Esther, discussed the commonness of teacher reliance. When asked about the atmosphere of her programming lessons, Esther simply stated “Chaotic …lots of hands up and down all the time.” and that these students’ problems are “almost always about errors in code.”. Chloe similarly reports “I think at least half the class will ask all the time.”; far beyond any teacher’s capacity to meaningfully assist every student.
Several teachers confidently posited that the cause of such reliance is a lack of resilience or motivation, rather than the difficulty of the error they are debugging. Some students want their errors to be instantly resolved, which they cannot always do themselves. This creates a preference for waiting for teacher support over sustained independent effort.
@@
Once a student raises their hand, teachers generally state that students will continue debugging until their teacher approaches them. This is of course negative for the student, with emotional reactions likely to ensue. However, this reliance also has a toll on the teacher. Some teachers reported that ‘programming lessons are really, really busy lessons’ (Chloe), which can “erode their [teachers’] professional confidence” (Adam). Esther talks about the obligation they feel to help every student, while Gideon feels the challenge of not being able to help every student.
@@
The impact of this pressure on the teacher is multi-faceted. As well as the ‘pressure to go and help everybody’, many teachers will feel a pressure to prove that their students have learnt something every lesson. This is very difficult when so many students’ progress depends on the individual support of the teacher.
Some teachers accept that it is not possible to personally help every student in need of debugging support. Instead, teacher reliance must be carefully managed. As Adam states, “there has to be a more scalable model than going around and solving individual people’s code problems”.
Consequently, strategies that motivate teacher-independent debugging have been utilised by some teachers. These aim to provide students with the confidence to debug without the teacher by their side. For example, some teachers tell their students to consult other sources of information before approaching them. These may be resources that teachers have curated (signposting to resources for assistance) or their peers (encouraging helpful collaboration). The latter does not just reduce the in-class burden on the teacher; explanations from peers “just seem to go in better” (Hannah). Alternatively, some teachers will simply refuse to help students until they have taken some effort to solve it themselves, such as Hannah.
@@
As a result, the teachers’ role regarding debugging support becomes less diagnostic and more flexible. In Isabel’s words, “they do ask [for debugging support], we just redirect them”. Students exposed to these strategies consequently become accustomed to teacher-independent debugging instead of teacher-dependent debugging. Teachers who employ such approaches explain how this helps to develop a resilience among students that is often lacking at the beginning of their programming journey.
‘They get better at these things, but there’s just as many older students who make the same mistakes and they’re like ‘oh, that was a silly error rather than ‘that was a killer.’ - Benjamin
@@ 4.4 Varying levels of scaffolding for helping students to debug
Although students can be redirected, there are of course times when teachers do need to provide debugging support to an individual student. This initiates a teacher-student interaction, where a teacher may provide a range of scaffolding techniques. The amount of scaffolding a teacher applies partially depends on teachers’ knowledge of the students, as well as practical factors such as lesson time remaining (Elliott et al., 2024).
Some scaffolding methods are generic teaching tactics applied to the context of debugging, while others are programming- or debugging-specific. The level of scaffolding varies across a continuum, from very little, placing the onus on the student, to near-complete scaffolding of the problem, placing the onus on the teacher.
When putting the onus on the student, some teachers simply ask students to share their thoughts about the problem at hand (encouraging externalisation of ideas) or ask students questions. Both of these approaches encourage students to outwardly reflect on their understanding of the problem, which might be what is stopping them from successfully resolving errors.
@@
The questions teachers mentioned were naturally of varying specificity, which again shifts the level of scaffolding. Adam deliberately uses generic questions when beginning interactions with his students.
@@
These questions are useful for specialist and non-specialist teachers, as they do not require expert knowledge nor time comprehending students’ code. Furthermore, once students understand they will be asked generic questions, they may be more inclined to debug more independently of the teacher.
Heavier forms of scaffolding may be used if a student has been trying to debug an error for a sustained period of time, or the error is difficult to localise. Some teachers may give students the solution or give students hints specific to the error(s) the student is struggling with. The latter is particularly useful when localising an error.
“If they’re struggling to find it, I can see they’re getting frustrated, first thing I would is I’d say, ‘Have a look at that particular line’. This is assuming I know where the error is.” - Benjamin

### v3

@@ 3.3 Analysis
We analysed teachers’ interviews using ‘codebook’ thematic analysis (TA) (ThematicAnalysisOriginalArticle; ConceptualDesignThinkingThematicAnalysis). This is a ‘school’ of TA that maintains the core values of reflexive TA, in which the subjectivity and experiences of the researcher are key aspects of the analysis, while incorporating some pragmatic approaches that seek reproducible coding (ConceptualDesignThinkingThematicAnalysis). In practice, we conceptualised latent themes that identified meaning beyond teachers’ responses, as discussed in ReflectingReflexiveThematicAnalysis; GoodPracticeThematicAnalysis, while incorporating approaches to ensure the codebook was consistently applied across the interview data (see Section 3.4).
This began with multiple passes of the interview transcripts. The first of these, as previously mentioned, was used both for correcting the automatically generated transcripts and initial memoing. This was performed by the first author. The second author then read through the corrected transcripts and wrote initial memos.
After this, the interview transcripts were migrated to NVivo 12, which was used for the rest of the analysis. Each transcript was then read through again by the first author to inductively generate open descriptive codes and note further memos. Through further inspection of the interview data, open codes were inductively abstracted into higher-level latent themes and sub-themes that aimed to provide an interpretive, and hence subjective, account of the data. This process of theme generation involved repeatedly restructuring the codebook, redefining the themes and sub-themes, and rewriting the report of the codebook, encompassing steps 4-6 in ThematicAnalysisOriginalArticle. This process was conducted by the first author with feedback from the second author after some of the iterations. Our approach was therefore more iterative than linear, which is considered a normal part of TA (ThematicAnalysisOriginalArticle).
@@ 3.4 Reliability and Validity
The content validity of the interview questions was strengthened through two procedures. First, the questions were iteratively refined by both authors to ensure they appropriately discussed teachers’ experiences, perspectives, and observations of debugging. Second, a pilot interview was conducted with a teacher to assess the wording of the questions and the timings of the interview. This teacher was personally contacted by the first author and received the same bookshop voucher as the other participating teachers. Their transcript was not included in the analysis.
The reliability with which the codebook was applied was improved by seeking agreement between the authors’ coding patterns without calculating an interrater reliability (IRR) coefficient. UnitisationProblemSemiStructuredInterviews call this negotiated agreement. We wanted to verify the consistency with which the first author coded the transcripts according to the codebook definitions, also known as intracoder reliability (IntercoderReliabilityQualAnalysisGuidelines). Seeking agreement by identifying disagreements in our coding was appropriate for this case (UnitisationProblemSemiStructuredInterviews; IRRHCI). This was particularly important as the first author was inexperienced in qualitative analysis. Negotiated agreement hence improved the clarity of definitions and the consistency with which the first author coded.
Calculating an IRR coefficient was not required for the above process and was considered detrimental for our study. Our interview transcripts were long and conversational, including frequent tangents and interruptions. Naturally, this makes it difficult for coders to agree on units of meaning, which is required for calculating agreement with IRR coefficients. Combined with the complexity of the codebook, any IRR coefficient is likely to be positively skewed and unrepresentative of the reliability of the codebook (UnitisationProblemSemiStructuredInterviews). Furthermore, there is little consensus on acceptable IRR coefficient values for semi-structured interview data, which undoubtedly differs from other qualitative data with different research objectives (ReliabilityQualAnalysisMisconceptions; UnitisationProblemSemiStructuredInterviews).
Agreement was hence sought as follows. As described in the previous section, the first author developed the codebook with regular input from the second author. Two interview transcripts (Gideon and Hannah) were then randomly selected to be coded by both authors, in line with the guidelines of IntercoderReliabilityQualAnalysisGuidelines. For this, the second author was provided with the codebook, complete with definitions, and an uncoded version of the transcripts. Once we both coded the transcripts, disagreements in coding were identified using the coding stripes feature in NVivo. Disagreements and ambiguous sub-themes were then discussed. Where necessary, revisions to the structure and definitions of the codebook were made.
@@ 3.5 Limitations
A limitation of our method was that we kept the final themes close to the research questions (i.e., closely related to debugging). This was done to avoid generating an overly large codebook, but in the process likely missed some information that contributes to teachers’ experiences of debugging in the classroom.
Additionally, our sample was, within certain constraints, self-selecting. Although we were not seeking a generalisable population for an in-depth qualitative study, the participating teachers were likely engaged with computing education research to some degree. As a result, they may be more reflective or intentional about their teaching practice than most teachers, potentially affecting the themes we identified.
@@ 4 Results
The codebook structure and code definitions are available to view in the study repository (StudyRepository). We now report each of the themes, with sub-themes reported in italics, supported by teachers’ quotes [^3] All teachers have been given pseudonyms and all quotes have had fillers and contractions removed. from the interviews.
@@ 4.1 Barriers to successful debugging
A particularly widespread theme was the obstacles that impede students from successfully debugging their programs. These challenges related to schools’ allotted time for teaching programming, students’ attitudes, and the nature of the Python language.
@@
Barriers related to the difficulties of learning to program (programming-related), some of which were specific to learning a TBPL, also make debugging difficult for students. The difficulty of getting syntax exactly right and conceptual difficulties, sometimes manifesting as logical errors, were both mentioned as struggles for students. Many teachers described the common errors students encounter, mostly related to minor incorrect formatting, illustrating the ‘syntax barrier’ (UnderstandingSyntaxBarrier) present for many beginner programmers. Esther makes the point that such syntax errors may not be minor ‘slips of the hand’, but a “a lack of the deeper understanding of what the language is actually trying to do”.
@@
Gideon highlights that students’ negative attitudes span from debugging as a skill to computing as a subject. He also alludes to a lack of success and motivation as drivers for these negative attitudes. The former relates to the ‘performance accomplishment’ determinant of self-efficacy (OriginalSelfEfficacyArticle), which is derived from a learner’s initial experiences with a subject. For most students in England, initial experiences of programming come at primary school in block-based environments. However, unhelpful attitudes could also arise from students’ experiences of learning more generally, which Isabel believes to be the case.
@@
The effect of these attitudes often manifests as a lack of motivation to debug their programs. This has consequences for the teacher, who has the challenge of encouraging their students to persevere with debugging.
@@ 4.2 The emotional nature of debugging
In addition to attitudes, teachers discussed a range of emotional reactions they observe when their students are debugging. Observed reactions were generally physical or verbal, occurring at the point of encountering an error or after attempting to resolve an error. Visible reactions are not universal, however. Some teachers specifically pointed out the reserved reactions of quieter students.
Most observed reactions represented negative emotion during debugging struggle, conveying feelings such as frustration and disappointment. An interesting observation mentioned by several teachers is the emotional attachment students develop towards their programs and the distress this can cause when errors occur. For students, encountering errors in their own programs can be far more costly than debugging an erroneous program written by the teacher, leading to more dire emotional reactions. Felix describes this as potentially depressing for students.
“When they’re starting to write upwards of ten lines of code, you’re invested in getting it right and then if it fails, that can flip and that can be quite depressing.” - Felix
@@
“Once they’ve hit the brick wall, that’s kind of the green light to go, ‘I’m just gonna do anything’. [They] Sit there in silence, they’re not causing any commotion.” - Gideon
Despite the lack of visible frustration, disengagement may indicate more severe emotional distress, as students may not be able to cope with any more errors in their programs. Alternatively, students may disengage due to the difficulty of the error. Either way, some students’ motivation to debug is reduced after repeatedly encountering errors. It can be difficult for busy teachers to notice these students who, as Gideon mentions, are “not causing any commotion”.
@@
Felix’s observations allude to the physical enactment of students’ emotions, demonstrating the sheer joy that can be experienced from debugging. However, such positive reactions were rarely mentioned, perhaps due to the perseverance needed to experience them.
We also found some teachers alluded to the resilience of motivated students, who do not experience such polarising emotions when debugging. These students remain composed and explore different solutions, not afraid to add errors to their programs. Hannah explains that most of her students have such resilience.
@@
The resilience to debug, to a large extent, comes from the student. From Hannah’s perspective, this derives from students’ interest in programming. Other factors are undoubtedly at play, such as the debugging-specific teaching strategies teachers use (see Section 4.5), their work in normalising errors in the classroom (see Section 4.6), and students’ access to computers at home. Hannah teaches in a highly selective school, where high levels of interest are perhaps not as surprising. Generally, however, negative emotions and disengagement were the most widely discussed observations we identified in this theme.
@@ 4.3 Reliance on the teacher
A combination of the barriers to successful debugging and the emotional nature of debugging contributes to many students relying on the teacher for debugging support. This is typically embodied by students raising their hands when encountering errors. Often, there are enough hands up to make it unrealistic for the teacher to effectively help every student requesting support.
Several teachers discussed the commonness of teacher reliance. When asked about the atmosphere of her programming lessons, Esther simply stated “chaotic …lots of hands up and down all the time” and that her students’ problems are “almost always about errors in code”. Chloe similarly reports “I think at least half the class will ask all the time”; far beyond any teacher’s capacity to meaningfully assist every student in classes of 25-30 students (ExpandingCSChallengesExperiences; UKAverageClassSizes).
Several teachers confidently posited that the cause of such reliance is a lack of student resilience or motivation, rather than the difficulty of the error they are debugging. Some students want their errors to be instantly resolved, which they cannot always do themselves. This creates a preference for waiting for teacher support over sustained independent effort.
@@
Once a student raises their hand, teachers generally stated that students do not continue debugging until their teacher approaches them. This is naturally detrimental for the student, with emotional reactions likely to ensue. However, this reliance also has a toll on the teacher. Some teachers reported that “programming lessons are really, really busy lessons” (Chloe), which can “erode their [teachers’] professional confidence” (Adam). Esther talks about the obligation they feel to help every student, while Gideon discusses the challenge of not being able to do so.
@@
The impact of this pressure on the teacher is multifaceted. As well as the “pressure to go and help everybody”, many teachers also feel a pressure to prove that their students have learnt something every lesson. This is very difficult when so many students’ progress depends on the teacher.
Some teachers accept that it is not possible to personally help every student in need of debugging support. Instead, teacher reliance must be carefully managed. As Adam states, “there has to be a more scalable model than going around and solving individual people’s code problems”. Consequently, some teachers utilise strategies that motivate teacher-independent debugging. These aim to provide students with the confidence to debug without the teacher by their side. For example, some teachers tell their students to consult other sources of information before approaching them. These may be their peers (encouraging helpful collaboration) or resources that teachers have curated (signposting to resources for assistance). The former does not just reduce the in-class burden on the teacher; explanations from peers “just seem to go in better” (Hannah). Alternatively, some teachers will simply refuse to help students until they have taken some effort to debug themselves, such as Hannah.
@@
As a result, the teachers’ role regarding debugging support becomes less diagnostic. In Isabel’s words, “they do ask [for debugging support], we just redirect them”. Students exposed to these strategies become accustomed to teacher-independent debugging instead of teacher-dependent debugging. Teachers who employ such approaches explained how this helps to develop resilience among students that is often lacking at the beginning of their programming journey.
‘They get better at these things, but there’s just as many older students who make the same mistakes and they’re like ‘oh, that was a silly error’ rather than ‘that was a killer’.” - Benjamin
@@ 4.4 Varying levels of scaffolding for helping students to debug
There are of course times when teachers do need to provide individualised debugging support to students. This initiates a teacher-student interaction, where teachers mentioned using a range of scaffolding techniques. Some of these are generic teaching tactics applied to the context of debugging, while others are programming- or debugging-specific. The level of scaffolding varies across a continuum, from very little, placing the onus on the student, to near-complete scaffolding of the problem, placing the onus on the teacher. The amount a teacher applies to a particular interaction partially depends on the teacher’s knowledge of their student(s), as well as practical factors such as remaining lesson time (TeacherSupportForDebuggingPhysicalComputing).
When putting the onus on the student, some teachers simply ask students to share their thoughts about the error in question (encouraging externalisation of ideas or asking students questions). Both of these approaches encourage students to outwardly reflect on their understanding of the problem, which might be what is stopping them from successfully resolving errors.
@@
The questions teachers mentioned asking students were naturally of varying specificity, which again shifts the level of scaffolding. Adam deliberately uses generic questions at the beginning of interactions with his students.
@@
These questions are useful for specialist and non-specialist teachers alike, as they do not require expert knowledge or time spent comprehending students’ code. Furthermore, once students understand that they will be asked generic questions, they may be more inclined to debug independently of the teacher.
Heavier forms of scaffolding may be used if a student has been debugging for a sustained period of time, or if the error is difficult to localise. Some teachers may give students the solution or give students hints specific to the error(s) the student is struggling with.
“If they’re struggling to find it, I can see they’re getting frustrated, first thing I would say is, ‘have a look at that particular line’. This is assuming I know where the error is.” - Benjamin